The Royal Fruit Company produces two types of fruit drinks. The first type is 65% pure fruit juice, and the second type is 90% pure fruit juice. The company is attempting to produce a fruit drink that contains 80% pure fruit juice. How many pints of each of the two existing types of drink must be used to make 120 pints of a mixture that is 80% pure fruit juice?
Answer by stanbon(75887)

The Royal Fruit Company produces two types of fruit drinks. The first type is 65% pure fruit juice, and the second type is 90% pure fruit juice. The company is attempting to produce a fruit drink that contains 80% pure fruit juice. How many pints of each of the two existing types of drink must be used to make 120 pints of a mixture that is 80% pure fruit juice?
-----
Quantity Eq:: s + n = 120 pints
Juice Eq:: 65S+90n = 80*120
--------------------------------
Modify for elimination:
65s + 65n = 65*120
65s + 90n = 80*120
-----
Subtract and solve for "n":
25n = 15*120
n = (3/5)120 = 72 pints (amt of 90% needed)
s = 120-72 = 48 pints (amt of 65% needed)
-------------
Cheers,
Stan H.
---------